40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 45     Entry time: Thu Nov 1 11:45:30 2007
Author: tobin 
Type: Configuration 
Category: IOO 
Subject: Mode cleaner drag-wiping 
Andrey, Bob, David, John Miller, Rana, Rob, Steve, Tobin

Yesterday we vented the vacuum enclosure and opened up the chamber containing MC1 & MC3 by removing the access connector between that chamber and the OMC chamber. Rana marked MC1's location with dogs and then slid the suspension horizontally to the table edge for easy drag-wiping access. The optic was thoroughly hosed-down with the dionizer, in part in an effort to remove dust from the cage and the top of the optic. Drag-wiping commenced with Rob squirting (using the 50 microliter syringe) and Tobin dragging (using half-sheets of Kodak lens tissue). We drag-wiped the optic many (~10) times, concentrating on the center but also chasing around various particles and a smudge on the periphery. There remains one tiny speck at about the 7:30 position, outside of the resonant spot area, that we could not dislodge with three wipes.

Today we drag-wiped MC3. First we slid MC1 back and then slid MC3 out to the edge of the table. We disconnected the OSEM cables in the process for accessibility, and MC1 is perched at an angle, resting on a dog. We did not blow MC3 with the deonizer, not wanting to blow particles from MC3 to the already-cleaned MC1. We drag-wiped MC3 only three times, all downward drags through the optic center, with Steve squirting and Tobin dragging. Some particles are still visible around the periphery, and there appears to be a small fiber lodged near the optic center on the reverse face.

Andrey and Steve have opened up MC2 in preparation for drag-wiping that optic after lunch.
ELOG V3.1.3-